the perimeter of a rectangle is 36 cm. the area is 80 cm squared. find the length and width.

i know the answers i just need how to get them. this is part of my area math packet. please help!

Answer by jorel1380(3719)

2L+2W=36
2L=36-2W
L=18-W
--------------
W(18-W)=80
18W-W2=80
W2-18W+80=0
(W-10)(W-8)=0
W=8 or 10
Choosing the width to be 8 cm., we get the length to be 10 cm..